Helium is contained in a 3 π‘š3 rigid volume at 60℃ and 100 kPa. Find the heat transfer needed to increase the pressure to 1000 kPa.


1
Expert's answer
2022-03-03T12:04:57-0500

"\nu=\frac{p_1V_1}{RT_1}=108.4~mol,"

"T_2=\frac{p_2T_1}{p_1}=3330~K,"

"Q=\frac 32\nu R\Delta T=4~MJ."